# IntactKV: Improving Large Language Model Quantization by Keeping Pivot Tokens Intact

### v1

@@ Appendix E Evaluation Details
MT-bench employs a GPT-4 model to score the generated content. In our experiments, we find that the score given by GPT-4 can vary for the same content even when the generation temperature of GPT-4 is set to 0. Besides, content generation for the writing and roleplay category has a relatively high generation temperature of 0.7, which also results in variations in the final score. To faithfully assess the performance of the quantized model, we run the content generation process of each model 3 times with random seeds 42, 43, and 44. We report the mean score of three trials as the final score in Table 5 and Table 6. Also, we note that GPT-4-Turbo has been shown to be smarter than GPT-4888https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard, and in our experiments, we find that GPT-4-Turbo can give more stable score than GPT-4. Therefore, we evaluate the generation results on MT-bench with the latest gpt-4-0125-preview API (i.e., GPT-4-Turbo) provided by OpenAI to further reduce variations in the final score.
Appendix F More Experiment Results
F.1 PPL Results on WikiText2
As shown in Table 7, AWQ+IntactKV consistently improves AWQ for every model on the WikiText2 dataset and outperforms OmniQuant for five out of seven models. We note that OmniQuant uses the WikiText2 dataset as a calibration set, while AWQ uses Pile dataset, which may give OmniQuant a bonus for PPL evaluation on the WikiText2 dataset.
F.2 MMLU Results
@@
We provide INT3-g128 quantization results on MMLU in Table 8, and INT4-g128 quantization results on MMLU in Table 9. For INT3-g128 quantization, AWQ+IntactKV consistently improves AWQ in every experiment setting and outperforms OmniQuant for nine out of ten settings. For INT4-g128 quantization, AWQ+IntactKV leads to relatively less improvement over AWQ on the MMLU benchmark compared with INT3-g128 quantization, but still outperforms AWQ in nine out of ten experiment settings, and performs on par with OmniQuant on MMLU.
F.3 Commonsense QA Results
@@
We conduct experiments on seven zero-shot commonsense QA tasks for the Vicuna family with both INT3-g128 and INT4-g128 quantization. The results are shown in Table 10. For INT3-g128 quantization, AWQ+IntactKV significantly surpasses all baselines in all experiment settings. For INT4-g128 quantization, AWQ+IntactKV consistently improves AWQ and outperforms OmniQuant in four out of five experiment settings, demonstrating the necessity for keeping an intact KV cache for pivot tokens.
Appendix G Experiment Details of KV Cache Quantization
We apply asymmetric per-head dynamic quantization to the KV cache. When combined with IntactKV, we still keep IntactKV in FP16 while the KV cache of other tokens in low bit, which only induces negligible memory overhead compared with quantizing the entire KV cache to low bits since IntactKV only contains the KV cache of a few tokens.
Appendix H Links to Officially Released LLMs

### v2

@@ Appendix F Evaluation Details
MT-bench employs a GPT-4 model to score the generated content. In our experiments, we find that the scores given by GPT-4 can vary for the same generated content even when the generation temperature of GPT-4 is set to 0. Besides, content generation for the writing and roleplay categories has a relatively high generation temperature of 0.7, which also results in variations in the final score. To faithfully assess the performance of the quantized model and decrease the variations in the final score, we run the content generation process of each model 3 times with random seeds 42, 43, and 44. We report the mean score of three trials as the final score in Table 5 and Table 13. Also, we note that GPT-4-Turbo has been shown to be smarter than GPT-4999https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard, and in our experiments, we find that GPT-4-Turbo can give more stable scores than GPT-4 while having a much lower price. Therefore, we evaluate the generation results on MT-bench with the latest gpt-4-0125-preview API (i.e., GPT-4-Turbo) provided by OpenAI to further reduce variations in the final score.
Appendix G More Experiment Results
G.1 PPL Results
We provide PPL results of LLaMA and LLaMA-2 models on WikiText2 in Table 7, and PPL results of LLaMA-3 models in Table 8. These results affirm IntactKV’s effectiveness in restoring the capabilities of quantized models. Moreover, in Table 9, we conduct experiments on more heterogeneous backbones like OPT and Mistral, which further proves the compatibility of our IntactKV with various LLM backbones.
G.2 MMLU Results
@@
We provide INT3-group128 weight-only quantization results on MMLU in Table 10, and INT4-group128 weight-only quantization results on MMLU in Table 11. For INT3-group128 quantization, AWQ+IntactKV consistently improves AWQ in every experiment setting and outperforms OmniQuant for nine out of ten settings. For INT4-group128 quantization, AWQ+IntactKV leads to relatively less improvement over AWQ compared with INT3-group128 quantization, but still outperforms AWQ in nine out of ten experiment settings, and performs on par with OmniQuant.
G.3 Commonsense QA Results
@@
We conduct experiments on seven zero-shot commonsense QA tasks for the Vicuna family with both INT3-group128 and INT4-group128 weight-only quantization. The results are shown in Table 12. For INT3-group128 quantization, AWQ+IntactKV significantly surpasses all baselines in four out of five experiment settings. For INT4-group128 quantization, AWQ+IntactKV improves AWQ and outperforms OmniQuant in four out of five experiment settings, demonstrating the superiority of IntactKV.
G.4 MT-Bench Results
We provide INT4-group128 quantization results on MT-bench in Table 13. As can be seen, IntactKV leads to an average increase of 0.09 in the final score. Remarkably, with trainable IntactKV, AWQ even matches the full-precision model under INT4 quantization, while all other methods clearly lag behind the full-precision model.
Appendix H Effectiveness of Calibrating IntactKV
We conduct more experiments on MT-Bench to further demonstrate the effectiveness of calibrating IntactKV. We adopt a commonly used fine-tuning method for quantized models that tunes the quantization bias term (used in non-symmetric quantization) in every quantization group as a baseline method, termed "group bias tuning". Both "group bias tuning" and "calibrating IntactKV" are further tuned based on "AWQ+IntactKV${}_{\textrm{[P]}}$". We use the same calibration set containing 128 samples and train 20 epochs for a fair comparison. As shown in Table 14, although calibrating IntactKV uses fewer trainable parameters, it still achieves better or comparable results compared with group bias tuning, demonstrating the effectiveness of calibrating IntactKV. Also, we note that "calibrating IntactKV" can be adopted for any quantization setting, while "group bias tuning" is only suitable for non-symmetric and group-wise quantization, making our proposed method a more versatile calibration strategy for quantized models.
Appendix I Adapting IntactKV for Activation Quantization
It is non-trivial to integrate IntactKV into activation quantization. For activation quantization, the whole KV cache needs to be quantized to low bits to exploit integer multiplications in self-attention, which contradicts our idea of keeping pivot tokens’ KV cache intact. However, as shown in Table 15, the distribution of the pivot tokens’ KV cache is much smoother than that of the non-pivot tokens’ KV cache, which implies that IntactKV is amenable to quantization. Therefore, we adopt a straightforward solution to adapt IntactKV for activation quantization that directly quantizes IntactKV to lower bits with RTN. As shown in Table 16, quantizing IntactKV incurs minimal accuracy loss. For example, quantizing IntactKV to 4 bits only results in an average PPL increase of 0.05 on WikiText2 compared with full-precision IntactKV, which is negligible.
Method INT3-group128 INT4-group128
AWQ 5.17 5.28
    +IntactKV${}_{\textrm{[P]}}$ 5.34 5.35
    +IntactKV${}_{\textrm{[P]}}$+gbias 5.31 5.47
    +IntactKV${}_{\textrm{[P]}}$+Cal 5.44 5.50
Table 14: Evaluation of different calibration methods on MT-bench. "gbias" denotes group bias tuning and "Cal" denotes calibrating IntactKV.
Method Pivot K Cache Pivot V Cache Non-pivot K Cache Non-pivot V Cache
AbsMax Std AbsMax Std AbsMax Std AbsMax Std
LLaMA-7B 3.15 0.38 0.63 0.04 13.91 1.58 2.34 0.46
LLaMA-13B 3.02 0.35 0.73 0.05 13.69 1.56 2.62 0.49
LLaMA-2-7B 2.76 0.30 0.79 0.05 14.28 1.65 2.23 0.42
LLaMA-2-13B 2.73 0.27 0.75 0.05 14.60 1.62 2.57 0.44
LLaMA-3-8B 3.30 0.37 0.57 0.03 15.86 2.19 1.54 0.27
Table 15: The statistical results of pivot tokens’ and non-pivot tokens’ KV cache. The maximum absolute value and standard deviation are calculated on a sequence of length 1024 and averaged over all layers.
Method LLaMA-7B LLaMA-13B LLaMA-2-7B LLaMA-2-13B
C4 WikiText2 C4 WikiText2 C4 WikiText2 C4 WikiText2
OmniQuant 17.03 12.17 15.65 11.16 21.40 14.74 16.24 12.28
    +IntactKV${}_{\textrm{[B]}}$ (FP16) 16.26 11.30 13.89 10.00 19.97 13.61 15.77 10.94
    +IntactKV${}_{\textrm{[B]}}$ 16.24 11.32 13.87 10.04 20.01 13.70 15.91 11.00
Table 16: The effect of quantizing IntactKV to lower bits. We show the INT4 weight and activation quantization results of LLaMA models on C4 and WikiText2 datasets. IntactKV${}_{\textrm{[B]}}$ (FP16) indicates keeping IntactKV in 16 bits, which incurs extra inference costs. IntactKV${}_{\textrm{[B]}}$ indicates quantizing IntactKV to lower bits (i.e., 4 bits).
Model Download URL
LLaMA-2-7B https://huggingface.co/meta-llama/Llama-2-7b
LLaMA-2-13B https://huggingface.co/meta-llama/Llama-2-13b
LLaMA-2-70B https://huggingface.co/meta-llama/Llama-2-70b
LLaMA-3-8B https://huggingface.co/meta-llama/Meta-Llama-3-8B
LLaMA-3-70B https://huggingface.co/meta-llama/Meta-Llama-3-70B
Vicuna-v1.3-7B https://huggingface.co/lmsys/vicuna-7b-v1.3
Vicuna-v1.3-13B https://huggingface.co/lmsys/vicuna-13b-v1.3
Vicuna-v1.3-33B https://huggingface.co/lmsys/vicuna-33b-v1.3
Vicuna-v1.5-7B https://huggingface.co/lmsys/vicuna-7b-v1.5
Vicuna-v1.5-13B https://huggingface.co/lmsys/vicuna-13b-v1.5
Table 17: Download links to officially released LLMs.
Appendix J Links to Officially Released LLMs
@@ Appendix I Adapting IntactKV for Activation Quantization
Figure 13: Magnitude of the output activations and attention map in LLaMA-3-8B.
Figure 14: Magnitude of the output activations and attention map in LLaMA-3-70B.
@@
Figure 20: Magnitude of the output activations and attention map in OPT-6.7B.
Figure 21: Magnitude of the output activations and attention map in Mistral-7B.